OpenStudy (anonymous):

5/16 - (-5/16). perform the indicated operation.

OpenStudy (anonymous):

\[\frac{5}{8} \]

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

I couldn't remember if I added the numerator.

OpenStudy (anonymous):

You end up adding 5/16 to 5/16 which is 10/16 or 5/8 because of the change of signs in the second fraction. \[\left(\frac{5}{16}-\left(\frac{-5}{16}\right)\right)=\left(\frac{5}{16}-\frac{-5}{16}\right)=\left(\frac{5}{16}+\frac{5}{16}\right)=\left(\frac{10}{16}\right)=\frac{5}{8} \]